# Adsorption of polyelectrolytes in the presence of varying dielectric discontinuity between solution and substrate

### v2

@@ Abstract
Polyelectrolytes (PEs), i.e., charged polymers, are of great interest in physics, materials science, chemistry, and biology. PE adsorption onto substrates is a particularly important process with applications covering, e.g., water purification, production of paper and cardboard, energy storage Zhu et al. (2021), chemical sensors Feng et al. (2010); Kim et al. (2011), stabilization of colloids Napper (1983); Eisenriegler (1993), and metal corrosion protection Andreeva, Skorb, and Shchukin (2010). It has been the subject of extensive experiments Maroni, Ruiz-Cabello, and Tiraferri (2014); Eneh et al. (2020); Ivanov et al. (2021); Li, Parakhonskiy, and Skirtach (2023); Li et al. (2023), theoretical investigations Netz and Joanny (1999); Dirir et al. (2015); Merlitz et al. (2015); Xie et al. (2016); Buyukdagli and Podgornik (2019), and particle-based simulations Luque-Caballero, Martín-Molina, and Quesada-Pérez (2014); De Carvalho, Metzler, and Cherstvy (2015); Kostritskii et al. (2016); Sánchez et al. (2019); Bagchi, Nguyen, and Olvera de la Cruz (2020); Wang et al. (2023). Biological examples of PEs are DNAs, RNAs, proteins, and microtubules. Their interactions with bio-based systems inevitably lead to interactions of the former with the cellular membranes, which are composed of a chemically diverse set of lipids Langecker et al. (2014). Additionally, PE assemblies are essential from the point of view of biomedical applications, including vesicle transport Di Paolo and De Camilli (2006); McLaughlin and Murray (2005), preventing bacterial adhesion Zhu et al. (2015); Ouni et al. (2021), and in drug and gene delivery systems Yin et al. (2014); Maier and Rädler (2000); Pott and Roux (2002); Pott et al. (2003); Kahl et al. (2009); Leal et al. (2008).
@@
As PE solutions and PE assemblies are in contact with various substrates in many of the applications, a thorough understanding of the fundamental PE-substrate interactions is needed. These interactions are highly complex and depend on substrate polarization, solution characteristics, relevant electrostatic forces stemming from PE line charge density, substrate surface charge density, co- and counterion valencies and image charges Bagchi, Nguyen, and Olvera de la Cruz (2020), steric and entropic effects (ion sizes, configurational entropies and PE flexibility Cherstvy and Winkler (2011, 2012); Dias and Pais (2012); Chang and Jiang (2021)), all of which influence the overall adsorption behavior and stability of the system Elizarova and Luckham (2018); Zhu et al. (2018). Straightforward electrostatic considerations show that for $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, the image charges induced on the substrate are repulsive, while for $\varepsilon_{\rm s}>\varepsilon_{\rm w}$ the response is opposite, namely attractive. To overcome the repulsive image charges, many works on PE-substrate adsorption have focused on the case of adding attractive charges on the substrate Dobrynin, Rubinstein, and Joanny (1997); Netz and Joanny (1999); Messina (2004); Cherstvy and Winkler (2011, 2012); Bagchi, Nguyen, and Olvera de la Cruz (2020); Wang et al. (2023) or applying an external electric field Son and Wang (2021). The situation for high-dielectric constant, metallic-character substrates is more subtle, however. Although the image charges are attractive, a competition between the interactions mentioned above exists Netz (2002); Cheng and Lai (2004); Seijo et al. (2009); Lee, Kostinski, and Brenner (2018). While adsorption of isolated PEs on metallic-character substrates due to attractive images is possible, the spontaneous formation of dense layers is a nontrivial issue due to the complex interactions.
@@
In this paper, we employ coarse-grained MD simulations to study PE solutions at electrostatically uncharged planar interfaces. We consider a simplified model of a rigid DNA-like molecule that experiences a dielectric contrast between the solution and an impenetrable substrate. We scrutinize the interactions and adsorption processes of a single and a pair of identical rodlike PEs with equal charge and address questions that arise regarding how substrate polarization can influence the adsorption phenomena, which are pronounced in the presence of multivalent ions in the salt. We focus on the effect of salt concentration, ion valency, and substrate characteristics, which are key aspects controlling adsorption as observed in experimental settings. For example, in the case of PE multilayers, one of the most influential variables on film thickness is the concentration of salt in the deposition solution Decher and Schmitt (1992); Dubas and Schlenoff (1999). We also anticipate that correlation effects will be significant, particularly in cases involving asymmetric salts that feature a low concentration of an ionic species with high valency Antila, Van Tassel, and Sammalkorpi (2017). The experimental relevance of this has been noted in previous studies Rau and Parsegian (1992a, b).
@@ II Simulation model
Due to the dielectric mismatch between the substrate and water, electric charges are induced on the substrate plane at $z=0$ in response to the presence of ions and PEs within the implicit water environment. We calculate the polarization charge density at each point on the substrate in response to a point charge $q$, as given by Jackson (1999)
@@
$\sigma_{\rm p}=-\frac{q\Delta{\varepsilon}}{2\pi\varepsilon_{\rm w}}\frac{d}{(z^{2}+d^{2})^{3/2}},$ (1)
where $\Delta{\varepsilon}=({{\varepsilon_{\rm s}-\varepsilon_{\rm w}}})/({{\varepsilon_{\rm s}+\varepsilon_{\rm w}}})$ is the dielectric mismatch between the substrate and solvent (water), $\varepsilon_{\rm 0}$ is the vacuum permittivity, $z$ is the vertical distance of the charge from the reference point on the substrate, and $d$ is the distance from the reference on the interface to the projection of the point charge on the interface. Here, the parameter $\Delta{\varepsilon}$ has values of $0.72$, $0.85$, and $0.98$ for $\varepsilon_{\rm s}$ equal to $500$, $1000$, and $8000$, respectively.
@@
Here, $Z$ represents the charge valency. The notation $Z^{\rm c}:Z^{\rm a}$ is used to represent the valencies of cations and anions, respectively. The Bjerrum length in water at $300$ K is $\ell_{\rm B}=\beta e^{2}/(4\pi\varepsilon_{0}\varepsilon_{\rm w})\approx 0.7$ nm, where $\beta=1/k_{\rm B}T$, with $k_{\rm B}$ being the Boltzmann constant and $T$ the temperature of the system. The Debye screening length, defined as $\kappa^{-1}=(4\pi\ell_{\rm B}\Sigma_{i}(Z^{i})^{2}c^{i})^{-1/2}$, represents the distance over which electrostatic interactions are exponentially screened by salt ions. For long-range electrostatic interactions beyond cutoff, we employ the Particle-Particle Particle-Mesh (P3M) summation procedure Plimpton, Pollock, and Stevens (1997) with slab corrections in the $z$ direction and a relative force accuracy of $10^{-5}$. $z$-periodicity is applied by adding empty space between the repetitions of the simulation box Yeh and Berkowitz (1999). The proportion of the extended dimension in the $z$ direction relative to the original dimension is set at $3$. A reflective wall is positioned at $z=15$ nm (or at $z=20$ nm for the PE perpendicular to the substrate). Periodic boundary conditions are applied in the $x$ and $y$ directions. All simulations are performed using LAMMPS 2 Aug 2023 stable release Plimpton (1995); Brown et al. (2009); Thompson et al. (2022) in the canonical ($NVT$) ensemble, employing the Nosé-Hoover Nosé (1984); Hoover (1985) thermostat to maintain $T=300$ K. The integration time step is set to $2$ fs, and the trajectories are sampled every $2$ ps over a $20$ ns simulation period, excluding the initial $5$ ns for equilibration. The initial configurations of the simulations are generated using Moltemplate Jewett et al. (2021) and Packmol Martínez et al. (2009).
We also incorporate polarizability due to the dielectric mismatch using the ICC (Induced Charge Computation) algorithm Tyagi et al. (2010), which has been incorporated into LAMMPS as detailed in Ref. Nguyen et al., 2019. To compute induced charges, we discretize the surface into patches. Each patch has an area of approximately $0.866\sigma^{2}$. The choice of this grid size strikes a balance between computational cost and accuracy. Simulating systems with dielectric discontinuities can be computationally demanding because of the complex many-body interactions associated with charge polarization at such interfaces. Employing a finer grid leads to significantly increased computational cost, whilst exceeding an area of $\sigma^{2}$ results in a noticeable reduction in computation accuracy.
@@ III.1 Single PE parallel to the substrate
To check the effect of image charges on the ion distributions, we first performed simulations for systems consisting exclusively of solvated ions without the presence of any PEs. Figure S1 of the Supplementary Material (SM) displays the number density profiles of cations and anions in a $1:-1$ salt solution with a concentration of $250$ mM. We observe that near the high dielectric constant substrate corresponding to a metallic character, both cations and anions have higher densities as compared to those close to low dielectric discontinuity substrate, which does not exhibit significant polarization. Conversely, near a low dielectric constant substrate (with insulating character), the ion densities are lower than those observed near a non-polarizable substrate Vertenstein and Ronis (1987); Frydel (2011). This phenomenon can be attributed to the interaction of ions with their image charges. When ions are close to a substrate bearing metallic characteristics, they encounter image charges of opposite signs, leading to attraction. In contrast, for a low dielectric constant, relatively insulating substrate, the image charges mirror the sign of the ions, resulting in repulsion.
@@ III.1.1 Influence of added monovalent salt
Next, we examine the influence of added salt on the interactions between a DNA-like PE and a substrate. Figure 2 illustrates the potential of mean force $V$ on the PE as a function of the distance between the PE and the substrate while varying both the salt concentration $Z^{\rm c}c^{\rm c}$ and the dielectric constant $\varepsilon_{\rm s}$. When $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, induced charges, resulting from the presence of the PE, bear the same charge as the PE. Consequently, the PE is repelled from the substrate, a phenomenon observed with DNA and neutral phospholipid membranes (typically having a relative dielectric permittivity of $2-4$), both experimentally Gromelski and Brezesinski (2006) and in MD simulations Antipina and Gurtovenko (2016). In contrast, when $\varepsilon_{\rm s}>\varepsilon_{\rm w}$, the induced charges change sign, and the PE is attracted to the substrate. In all cases, increasing salt concentration results in more effective screening of the image charge effects, thus reducing the interaction strength between the PE and the substrate. This is consistent with the experimental findings reported in Ref. Boulmedais et al., 2006, which demonstrated that increasing the salt concentration leads to a reduction in the adsorption of poly(L-lysine) and heparin on conductive indium tin oxide substrate.
Figure 2: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ from the center of the PE to the substrate plane at $z=0$ for salts with valency ratios of $1:-1$ (top row) and $3:-1$ (bottom row), while varying both the salt concentration $Z^{\rm c}c^{\rm c}$ and the relative dielectric constant of the substrate $\varepsilon_{\rm s}$ (see Fig. S2 of SM for $2:-1$ salt). The vertical dashed lines represent the distance at which the surface of the PE is in contact with the substrate beads $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart such that $z_{2}=z_{1}+\sigma^{i}$.
@@ III.1.2 Influence of multivalent counterions
We next focus on the influence of multivalent ions on PE-substrate interactions. Incorporating multivalent cations into the electrolyte significantly reduces PE-substrate electrostatic interactions. This implies that higher-charged cations are more efficient in screening electrostatic interactions between the PE and the substrate. For example, for $\varepsilon_{\rm s}=8000$, the introduction of trivalent cations results in a weaker attraction between the PE and the substrate than that observed with monovalent cations, as evidenced by comparing Figs. 2(a) and 2(i). This observation is in contrast with the interactions between two nonpolarizable charged entities, such as two PEs or a PE with a substrate, where multivalent cations trigger charge reversal, leading to an attraction between the entities (see, e.g., Refs. Grosberg, Nguyen, and Shklovskii, 2002; Moreira and Netz, 2000; Buyukdagli and Ala-Nissila, 2017; Vahid et al., 2023 for a detailed discussion). Additionally, in $1:Z^{\rm a}$ salts we note that an increase in anion valency does not substantially alter the interactions between the PE and the substrate, as demonstrated in Fig. S3 of the SM.
@@
The results presented in Fig. 2(f) indicate that the potential curves for both high dielectric constant, metallic-character substrates, and those that have low dielectric constants tend to converge towards those of a non-polarizable substrate ($\varepsilon_{\rm s}=80$) with highly multivalent cations and at the highest salt concentrations studied here. This further implies that the screening of electrostatic interactions between the substrate and the PE is significantly more effective with high salt concentrations and cation valencies. In all cases, $V$ is closer to those with $\varepsilon_{\rm s}=80$ when $\varepsilon_{\rm s}<\varepsilon_{\rm w}$ than when $\varepsilon_{\rm s}>\varepsilon_{\rm w}$. This indicates that substrates with a low dielectric constant have a less pronounced effect on altering interactions between the PE and the substrate as compared to high dielectric constant substrates that have metallic character in their response. Additionally, Fig. 2(f) illustrates that the potential of mean force for $\varepsilon_{\rm s}=8000$ approaches zero at separations $z\approx 1.5$ nm, which corresponds to less than one ion diameter from the substrate. Here, dense packing of multivalent counterions around the PE results in induced charges on the substrate that effectively neutralize those caused by the PE proximity.
@@
Figure 3: The average electrostatic potential $\phi$ (cf. Eq. (5)) as a function of radial distance $r$ from the PE center in (a) $1:-1$ and (b) $3:-1$ salt solutions with $Z^{\rm c}c^{\rm c}=500$ mM. Here, the PE is at $z=1.4$ nm. The gray bar indicates the PE with a radius of $1$ nm. (c)-(e) The distribution of the total two-dimensional charge density, $n_{\rm t}=n^{\rm c}+n^{\rm a}$, in $3:-1$ salt at $Z^{\rm c}c^{\rm c}=500$ mM for different $\varepsilon_{\rm s}$. The data are averaged along the $y$ axis. The center of the PE axis is at $x=0$ and $z=1.4$ nm, with alignment along the $y$ axis. The PEs are shown in gray.
@@
In Fig. 3, we present the profiles of the mean electrostatic potential $\phi$ at $Z^{\rm c}:-1$ salt when the PE is at $z=1.4$ nm from the substrate plane. This value corresponds approximately to the position of the global minimum in the potentials of Fig. 2 (where such minima exist). In the case of $1:-1$ salt, charge reversal and the associated reversal of $\phi$ are absent for the PE near high dielectric constant or non-polarizable substrates. However, a weak inversion of $\phi$ is observed when the PE is close to a low dielectric constant, insulating substrate with a dielectric constant of $\varepsilon_{\rm s}=2$, as the induced charges on the substrate, which have the same sign as the PE, also attract cations.

### v3

@@ Abstract
Polyelectrolytes (PEs), i.e., charged polymers, are of great interest in physics, materials science, chemistry, and biology. PE adsorption onto substrates is a particularly important process with applications covering, e.g., water purification, production of paper and cardboard, energy storage Zhu et al. (2021), chemical sensors Feng et al. (2010); Kim et al. (2011), stabilization of colloids Napper (1983); Eisenriegler (1993), and metal corrosion protection Andreeva, Skorb, and Shchukin (2010). It has been the subject of extensive experiments Maroni, Ruiz-Cabello, and Tiraferri (2014); Eneh et al. (2020); Ivanov et al. (2021); Li, Parakhonskiy, and Skirtach (2023); Li et al. (2023), theoretical investigations Netz and Joanny (1999a); Dirir et al. (2015); Merlitz et al. (2015); Xie et al. (2016); Buyukdagli and Podgornik (2019a), and particle-based simulations Luque-Caballero, Martín-Molina, and Quesada-Pérez (2014); De Carvalho, Metzler, and Cherstvy (2015); Kostritskii et al. (2016); Sánchez et al. (2019); Bagchi, Nguyen, and Olvera de la Cruz (2020); Wang et al. (2023). Biological examples of PEs are DNAs, RNAs, proteins, and microtubules. Their interactions with bio-based systems inevitably lead to interactions of the former with the cellular membranes, which are composed of a chemically diverse set of lipids Langecker et al. (2014). Additionally, PE assemblies are essential from the point of view of biomedical applications, including vesicle transport Di Paolo and De Camilli (2006); McLaughlin and Murray (2005), preventing bacterial adhesion Zhu et al. (2015); Ouni et al. (2021), and in drug and gene delivery systems Yin et al. (2014); Maier and Rädler (2000); Pott and Roux (2002); Pott et al. (2003); Kahl et al. (2009); Leal et al. (2008).
@@
As PE solutions and PE assemblies are in contact with various substrates in many of the applications, a thorough understanding of the fundamental PE-substrate interactions is needed. These interactions are highly complex and depend on substrate polarization, solution characteristics, relevant electrostatic forces stemming from PE linear charge density, substrate surface charge density, co- and counterion valencies and image charges Bagchi, Nguyen, and Olvera de la Cruz (2020), steric and entropic effects (ion sizes, configurational entropies and PE flexibility Cherstvy and Winkler (2011, 2012); Dias and Pais (2012); Chang and Jiang (2021)), all of which influence the overall adsorption behavior and stability of the system Elizarova and Luckham (2018); Zhu et al. (2018). Straightforward electrostatic considerations show that for $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, the image charges induced on the substrate are repulsive, while for $\varepsilon_{\rm s}>\varepsilon_{\rm w}$ the response is opposite, namely attractive. To overcome the repulsive image charges, many works on PE-substrate adsorption have focused on the case of adding attractive charges on the substrate Dobrynin, Rubinstein, and Joanny (1997); Netz and Joanny (1999a); Messina (2004); Cherstvy and Winkler (2011, 2012); Bagchi, Nguyen, and Olvera de la Cruz (2020); Wang et al. (2023) or applying an external electric field Son and Wang (2021). The situation for high-dielectric constant, metallic-character substrates is more subtle, however. Although the image charges are attractive, a competition between the interactions mentioned above exists Netz (2002); Cheng and Lai (2004); Seijo et al. (2009); Lee, Kostinski, and Brenner (2018). While adsorption of isolated PEs on metallic-character substrates due to attractive images is possible, the spontaneous formation of dense layers is a nontrivial issue due to the complex interactions.
@@
In this paper, we employ coarse-grained MD simulations to study PE solutions at electrostatically uncharged planar interfaces. We consider a simplified model of a rigid double-stranded B-DNA-like molecule Cherstvy (2005) that experiences a dielectric contrast between the solution and an impenetrable substrate. We scrutinize the interactions and adsorption processes of a single and a pair of identical rodlike PEs with equal charge and address questions that arise regarding how substrate polarization can influence the adsorption phenomena, which are pronounced in the presence of multivalent ions in the salt. We focus on the effect of salt concentration, ion valency, and substrate characteristics, which are key aspects controlling adsorption as observed in experimental settings. For example, in the case of PE multilayers, one of the most influential variables on film thickness is the concentration of salt in the deposition solution Decher and Schmitt (1992); Dubas and Schlenoff (1999). We also anticipate that correlation effects will be significant, particularly in cases involving asymmetric salts that feature a low concentration of an ionic species with high valency Antila, Van Tassel, and Sammalkorpi (2017). The experimental relevance of this has been noted in previous studies Rau and Parsegian (1992a, b).
@@ II Simulation model
Due to the dielectric mismatch between the substrate and water, electric charges are induced on the substrate plane at $z=0$ in response to the presence of ions and PEs within the implicit water environment. The polarization charge density at each point on the substrate in response to the electric field at that point $\bm{E^{i}}$, generated by the PE solution and induced charges on other points on the substrate surface, is given by Jackson (1999)
$\sigma_{\rm p}^{i}=2\varepsilon_{0}\Delta{\varepsilon}~{}\bm{n}\cdot\bm{E^{i}},\hskip 30.0pti\in{\rm s},$ (1)
where $\Delta{\varepsilon}=({{\varepsilon_{\rm s}-\varepsilon_{\rm w}}})/({{\varepsilon_{\rm s}+\varepsilon_{\rm w}}})$ is the dielectric mismatch between the substrate and the implicit solvent (water), $\varepsilon_{\rm 0}$ is the vacuum permittivity, and $\bm{n}$ is the substrate normal vector toward the solvent. Here, the parameter $\Delta{\varepsilon}$ has values of $-0.95$, 00, $0.72$, $0.85$, and $0.98$ for $\varepsilon_{\rm s}$ equal to $2$, $80$, $500$, $1000$, and $8000$, respectively.
@@
Here, $Z$ represents the charge valency. The notation $Z^{\rm c}:|Z^{\rm a}|$ is used to represent the valencies of cations and anions, respectively. The Bjerrum length in water at $300$ K is $\ell_{\rm B}=\beta e^{2}/(4\pi\varepsilon_{0}\varepsilon_{\rm w})\approx 0.7$ nm, where $\beta=1/k_{\rm B}T$, with $k_{\rm B}$ being the Boltzmann constant and $T$ the temperature of the system. The Debye screening length, defined as $\kappa^{-1}=(4\pi\ell_{\rm B}\Sigma_{i}(Z^{i})^{2}c^{i})^{-1/2}$, represents the distance over which electrostatic interactions are exponentially screened by salt ions. For long-range electrostatic interactions, we employ the Particle-Particle Particle-Mesh (P3M) summation procedure Plimpton, Pollock, and Stevens (1997) with slab corrections in the $z$ direction and a relative force accuracy of $10^{-5}$. $z$-periodicity is applied by adding empty space between the repetitions of the simulation box Yeh and Berkowitz (1999). The proportion of the extended dimension in the $z$ direction relative to the original dimension is set at $3$. A reflective wall is positioned at $z=15$ nm (or at $z=20$ nm for the PE perpendicular to the substrate). Periodic boundary conditions are applied in the $x$ and $y$ directions. All simulations are performed using LAMMPS 2 Aug 2023 stable release Plimpton (1995); Brown et al. (2009); Thompson et al. (2022) in the canonical ($NVT$) ensemble, employing the Nosé-Hoover Nosé (1984); Hoover (1985) thermostat to maintain $T=300$ K. The integration time step is set to $2$ fs, and the trajectories are sampled every $2$ ps over a $20$ ns simulation period, excluding the initial $5$ ns for equilibration. The initial configurations of the simulations are generated using Moltemplate Jewett et al. (2021) and Packmol Martínez et al. (2009).
We also incorporate polarizability due to the dielectric mismatch using the ICC (Induced Charge Computation) algorithm Tyagi et al. (2010), which has been implemented into LAMMPS as detailed in Ref. Nguyen et al., 2019. In the algorithm, the substrate surface ${\rm s}$ is discretized into surface elements. A charged particle $i$ interacts with ${\rm s}$ via the surface element charges that capture the effect of the image charges due to the particle itself and all other charges in the system, including all the ions and the PE charges. The resulting coupled equations are solved self-consistently.
@@ III.1 Single PE parallel to the substrate
To check the effect of image charges on the ion distributions, we first performed simulations for systems consisting exclusively of solvated ions without the presence of any PEs. Figure S1 of the Supplementary Material (SM) displays the number density profiles of cations and anions in a $1:1$ salt solution with a concentration of $250$ mM. We observe that near the high dielectric constant substrate corresponding to a metallic character, both cations and anions have higher densities as compared to those close to low dielectric discontinuity substrate, which does not exhibit significant polarization. Conversely, near a low dielectric constant substrate (with insulating character), the ion densities are lower than those observed near a non-polarizable substrate Vertenstein and Ronis (1987); Frydel (2011). This phenomenon can be attributed to the interaction of ions with their image charges. When ions are close to a substrate bearing metallic characteristics, they encounter image charges of opposite signs, leading to attraction. In contrast, for a low dielectric constant, a relatively insulating substrate, the image charges mirror the sign of the ions, resulting in repulsion.
@@ III.1.1 Influence of added monovalent salt
Next, we examine the influence of added salt on the interactions between a DNA-like PE and a substrate. Figure 2 illustrates the potential of mean force $V$ on the PE as a function of the distance between the PE and the substrate while varying both the salt concentration $Z^{\rm c}c^{\rm c}$ and the dielectric constant $\varepsilon_{\rm s}$. When $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, induced charges, resulting from the presence of the PE, bear the same charge as the PE. Consequently, the PE is repelled from the substrate, a phenomenon observed with DNA and neutral phospholipid membranes (typically having a relative dielectric permittivity of $2-4$), both experimentally Gromelski and Brezesinski (2006) and in MD simulations Antipina and Gurtovenko (2016). This implies that in PE adsorption onto low dielectric constant substrates, the substrate must carry static charges Cherstvy and Winkler (2012). Theoretical studies based on a one-loop expansion of the self-consistent equation and incorporating charge correlations have revealed that strongly charged PEs are attracted to highly like-charged substrates due to the charge correlations Buyukdagli and Blossey (2016); Buyukdagli and Podgornik (2019a). However, increasing the bulk salt concentration leads to the unbinding of the PE from the substrate Buyukdagli and Blossey (2016). Weakly charged PEs are repelled by like-charged substrates with $\varepsilon_{\rm s}<\varepsilon_{\rm w}$, irrespective of salt concentration and substrate static charge density Buyukdagli and Blossey (2016). Similarly, mean-field-based investigations have shown that strongly charged semiflexible PEs desorb from oppositely charged planar substrates with $\varepsilon_{\rm s}<\varepsilon_{\rm w}$ at low and high monovalent salt concentrations due to image charge and entropic effects, respectively Netz and Joanny (1999a). On weakly charged substrates, repulsive interactions originating from image charge effects keep strongly charged PEs away from insulating character substrates Buyukdagli and Blossey (2016). In contrast, when $\varepsilon_{\rm s}>\varepsilon_{\rm w}$, the induced charges change sign, and the PE is attracted to the substrate. In all cases, increasing salt concentration results in more effective screening of the image charge effects, thus reducing the interaction strength between the PE and the substrate. This is consistent with the experimental findings reported in Ref. Boulmedais et al., 2006, which demonstrated that increasing the salt concentration leads to a reduction in the adsorption of poly(L-lysine) and heparin on conductive indium tin oxide substrate. It has been shown, by means of Debye-Hückel theory, that at high concentrations of monovalent salt, semiflexible PEs partially or fully detach from oppositely charged spheres Netz and Joanny (1999b); Boroudjerdi and Netz (2005) and planar substrates Netz and Joanny (1999a) due to the screening of electrostatic interactions. At approximately $1$ M salt concentration, the PE and the sphere fully separate and reach equilibrium Netz and Joanny (1999b). However, PEs form a lamellar structure on the substrate at intermediate salt concentrations, leading to substrate overcharging Netz and Joanny (1999a, b).
Figure 2: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ from the center of the PE to the substrate plane at $z=0$ for salts with valency ratios of (a)-(c) $1:1$ and (d)-(f) $3:1$, while varying both the salt concentration $Z^{\rm c}c^{\rm c}$ and the relative dielectric constant of the substrate $\varepsilon_{\rm s}$ (see Fig. S2 of SM for $2:1$ salt). The vertical dashed lines represent the distance at which the surface of the PE is in contact with the substrate beads $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart such that $z_{2}=z_{1}+\sigma^{i}$.
@@ III.1.2 Influence of multivalent counterions
We next focus on the influence of multivalent ions on PE-substrate interactions. Incorporating multivalent cations into the electrolyte significantly reduces PE-substrate electrostatic interactions. This implies that higher-charged cations are more efficient in screening electrostatic interactions between the PE and the substrate. For example, for $\varepsilon_{\rm s}=8000$, the introduction of trivalent cations results in a weaker attraction between the PE and the substrate than that observed with monovalent cations, as evidenced by comparing Figs. 2(a) and 2(i). Our findings are consistent with the experimental observations for DNA-cationic substrates, where adding multivalent cations releases DNA from the substrate Koltover, Wagner, and Safinya (2000). Additionally, in $1:|Z^{\rm a}|$ salts, we note that an increase in anion valency does not substantially alter the interactions between the PE and the substrate, as demonstrated in Fig. S3 of the SM.
On the other hand, one might expect to have like-charge attraction between the PE and the substrate with $\varepsilon_{\rm s}<\varepsilon_{\rm w}$ in solutions with multivalent ions, see e.g. Refs. Grosberg, Nguyen, and Shklovskii, 2002; Moreira and Netz, 2000; Buyukdagli and Ala-Nissila, 2017; Vahid et al., 2023 for detailed discussion. Theoretical studies, such as those using a one-loop expansion of self-consistent theory, have demonstrated that highly charged PEs experience stronger adsorption onto insulating substrates with like static charges in the intermediate coupling regime Buyukdagli and Podgornik (2019a). Furthermore, Ref. Buyukdagli and Podgornik, 2019b introduced a one-loop-dressed strong coupling theory to examine PE adsorption onto like-charged substrates in solutions containing tri- and tetravalent counterions with monovalent salt. The findings revealed that increasing counterion valency can enhance the adsorption of PEs, even on weakly charged substrates. Note that in these studies, the substrates have fixed charges that remain unaffected by the presence of ions and PEs. However, Fig. 2 shows that the PE attraction to the substrate with $\varepsilon_{\rm s}=8000$ is weakened by introducing multivalent cations to the salt. In our simulations, the bare charge density of the substrate is always zero, and the introduction of multivalent ions reduces, and in some cases completely neutralizes, the induced charges on the substrate. Consequently, the PE-substrate attraction significantly decreases. At high enough concentrations ($Z^{\rm c}c^{\rm c}\geq 60$ mM), the attraction completely vanishes.
@@
Figure 3: The average electrostatic potential $\phi$ (cf. Eq. (5)) as a function of radial distance $r$ from the PE center in (a) $1:1$ and (b) $3:1$ salt solutions with $Z^{\rm c}c^{\rm c}=500$ mM. Here, the PE is at $z=1.4$ nm. The gray bar indicates the PE with a radius of $1$ nm. (c)-(e) The distribution of the total two-dimensional charge density, $n_{\rm t}=n^{\rm c}+n^{\rm a}$, in $3:1$ salt at $Z^{\rm c}c^{\rm c}=500$ mM for different $\varepsilon_{\rm s}$. The data are averaged along the $y$ axis. The center of the PE axis is at $x=0$ and $z=1.4$ nm, with alignment along the $y$ axis. The PEs are shown in gray.
The results presented in Fig. 2(f) indicate that the potential curves for both high dielectric constant, metallic-character substrates, and those that have low dielectric constants tend to converge towards those of a non-polarizable substrate ($\varepsilon_{\rm s}=80$) with highly multivalent cations and at the highest salt concentrations studied here. This further implies that the screening of electrostatic interactions between the substrate and the PE is significantly more effective with high salt concentrations and cation valencies. In all cases, $V$ is closer to those with $\varepsilon_{\rm s}=80$ when $\varepsilon_{\rm s}<\varepsilon_{\rm w}$ than when $\varepsilon_{\rm s}>\varepsilon_{\rm w}$. This indicates that substrates with a low dielectric constant have a less pronounced effect on altering interactions between the PE and the substrate as compared to high dielectric constant substrates that have metallic character in their response. Additionally, Fig. 2(f) illustrates that the potential of mean force for $\varepsilon_{\rm s}=8000$ approaches zero at separations $z\approx 1.5$ nm, which corresponds to less than one ion diameter from the substrate. Here, dense packing of multivalent counterions around the PE results in induced charges on the substrate that effectively neutralize those caused by the PE proximity. Additionally, a possible explanation for the lower steric barrier height in $3:1$ salt as compared to $1:1$ salt across all values of $Z^{\rm c}c^{\rm c}$ when $\varepsilon_{\rm s}=8000$ could be attributed to the significantly lower number of ions present in the high-valency salt.
@@
In Fig. 3, we present the profiles of the mean electrostatic potential $\phi$ at $Z^{\rm c}:1$ salt when the PE is at $z=1.4$ nm from the substrate plane. This value corresponds approximately to the position of the global minimum in the potentials of Fig. 2 (where such minima exist). In the case of $1:1$ salt, charge reversal and the associated reversal of $\phi$ are absent for the PE near high dielectric constant or non-polarizable substrates. However, a weak inversion of $\phi$ is observed when the PE is close to a low dielectric constant, insulating substrate with a dielectric constant of $\varepsilon_{\rm s}=2$, as the induced charges on the substrate, which have the same sign as the PE, also attract cations.